# Bumblebee gravity: spherically-symmetric solutions away from the potential minimum

### v1

@@ 1 $a_{f1}$ $72e^{4}{\tilde{g}}a_{f1}^{2}/\beta_{3}^{2}$ $-\beta_{3}/6e^{2}$
Now we consider the first hypergeometric potential $M(3,2,X)$. The behavior indicated by the sample in figure 6(c) suggests a regular solution for the functions $f$, $M$, and $N$ until one reaches $x=4.38$, where the numerical code fails (for the given initial conditions). We have analyzed this region (near $x=4.38$) with calculation of curvature scalars and we find disagreement (see the curvature section below). For example, the Ricci Scalar we find must be finite at $x=4.38$, in disagreement with the numerical calculations. This behavior is consistent when one varies the initial values of $f^{\prime}$ and the coupling ${\tilde{g}}$; different choices merely move the stiff point to larger or shorter $x$. Thus it remains an open question to decipher the behavior in these points, whether they are coordinate singularities or physical ones.
@@
For the case of $M(4,2,X)$, the results of the sample in figure 6(d) show an interesting pulse-like behavior similar to the quadratic potential case. Around $x=25$, the function $N$ hits zero again, indicated another possible horizon. Indeed, calculations of the curvature show that it is finite at $x=25$. Thus we can say that this case contains RN-like behavior with two horizons.
@@
Not shown is the case for $M(5,3,X)$, for which we find that the behavior is similar to the case $M(4,3,X)$. The calculation of the curvature also confirms that when $N$ approches zero, the curvature is finite as indicated from curvature invariants. For both $n=4,5$ hypergeometric potentials it remains an open problem to map out the spacetime beyond this value of $x$ for which one gets horizon behavior.
@@ V.1.2 Asymptotic region
We then solve order by order in $u$. Note that this implies that the coefficients in the expansion must take values to cancel terms with inverse powers of $u$, starting with $1/u^{4}$. We include tables with the first 3 terms in the series for the different choices of potential.
@@ 1 $c_{f1}$ 00 $c_{N1}$ $c_{f1}$ 00 $2c_{f0}c_{f1}$
0 $c_{f0}$ $-c_{f0}/\sqrt{3}$ $c_{f0}^{2}/3$ $c_{f0}$ $-c_{f0}/\sqrt{2}$ $c_{f0}^{2}/2$
1 $c_{f1}$ 00 $\tfrac{2}{3}c_{f0}c_{f1}$ $c_{f1}$ 00 $c_{f0}c_{f1}$
2 $-c_{f1}^{2}/c_{f0}$ $-\frac{\sqrt{3}kc_{f1}^{2}}{2c_{f0}}$ 00 $-c_{f1}^{2}/c_{f0}$ $kc_{f1}^{2}/(\sqrt{2}c_{f0})$ 00
@@ 1 $c_{f1}$ 00 $\tfrac{2}{3}c_{f0}c_{f1}$ $c_{f1}$ 00 $c_{f0}c_{f1}$
The tables 4 and 5 display some general features for how the functions behave in the asymptotic region. Firstly, the initial data needed to determine the behavior rely only on $f$ and $df/du$ at $u=0$. For the non-RN cases of the quadratic and hypergeometric potentials, the signs of $c_{f0}$ and $c_{f1}$ are critical to the behavior of the solutions for increasing $u$, and thus decreasing $x$. Regarding the function $N$, if $c_{f1}c_{f0}>0$ then $N$ will be increasing up to order $u^{3}$, and the order $u^{2}$ term vanishes. This behavior resembles that of a naked singularity. Indeed this is born out by studying numerical solutions. On the other hand if $c_{f1}c_{f0}<0$, then $N$ will start off decreasing. For the latter condition, this is a good starting constraint on the initial values in order to try to find solutions that approach a horizon, like the conventional RN case.
The function $M$ has a different behavior. The $O(u)$ term vanishes for all potentials. The second and third order terms can be positive or negative depending on choices of $c_{f0}$ and $c_{f1}$. This was to be expected given the $M^{\prime}\sim M^{3}$ structure in (71).
@@
We can also confirm the Bekenstein result for the massive vector field, as the exact solution for this case emerges from the asymptotic expansion. Specifically, the solutions for the massive potential requires that $c_{f0}=0=c_{f1}$, so this case reduces to the Schwarzschild solution with zero vector field.
@@
(a) The quadratic potential case with asymptotic values for the functions of order unity. The initial values are taken to be $f(0)=-1$, $f^{\prime}(0)=1/2$ and we set ${\tilde{\lambda}}=1$.
(b) A second case of the quadratic potential with initial values of $f(0)=-1$, $f^{\prime}(0)=-1/5$, and ${\tilde{\lambda}}=1$. This behavior is consistent with a naked singularity.
(c) The hypergeometric potential $M(3,2,X)$ case. The asymptotic values used are $f(0)=-\sqrt{3}$ and $f^{\prime}(0)=1/\sqrt{3}$ while ${\tilde{g}}=1$ and $k=1/2$.
(d) The hypergeometric potential $M(4,2,X)$ case. The initial values are $f(0)=-\sqrt{2}$ and $f^{\prime}(0)=1/\sqrt{2}$, while $k=2/3$ and ${\tilde{g}}=1$.
Figure 7: Numerical solutions using the asymptotic expansion (70) to generate initial values near $u=0$. The analytical results with the series expansion are labeled with subscript $A$. The dashed lines are the RN case while the dotted lines show the analytical series from (70); the latter includes terms up to $u^{6}$. The connection with conventional Schwarzschild coordinates is $g_{tt}=A=-N$
@@
For the quadratic potential, shown in Figs. 7(a) and 7(b), two choices of initial values are made, one which yields decreasing behavior of $N$, the other increasing behavior. For the former, while $N$ appears headed to zero, $M$ increases without limit, in what appears to be a singular or stiff point in the numerical simulation. From curvature analysis, the point $u\approx 0.947$ appears to be a physical singularity of a different type from conventional ones (because $N\rightarrow 0$ and $M\rightarrow\infty$). We hesitate to call it a naked singularity because it may be partially hidden since $N\rightarrow 0$. For the latter case in 7(b), qualitatively different boundary conditions yield a clear naked singularity.
@@
For the $M(3,2,X/\Lambda^{2})$ case, we display a plot for which the asymptotic values will yield $dN/du<0$ initially. As shown in Fig. 7(c), the solution indeed appears to have $N$ reach 00. However, at this point the numerical code breaks down. Studies of curvature scalars show a smoothly increasing Kretschmann curvature scalar, revealing that there appears to be a singularity at $u\approx 1.90$. The smoothness gives an indicator of the reliability of the numerical solver [120]. This is similar to the case encountered for the quadratic potential.
For the $M(4,2,X/\Lambda^{2})$ case, we again display a plot for which the asymptotic values will yield $dN/du<0$ initially. As shown in Fig. 7(d), the solution appears to have $N$ reach 00. At this point, the numerical code breaks down; the curvature scalars show a smoothly increasing Kretschmann scalar, again indicating a singularity at $u\approx 1.12$.
As in the horizon expansion case in V.1.1, it would be of interest to further study these cases over a range of values for the initial data. We leave this for future work.
@@ V.1.3 Curvature
Invariant quantities can show unequivocally if singularities exist in the solutions. We first examine the Kretschmann scalar $K=R_{\mu\nu\kappa\lambda}R^{\mu\nu\kappa\lambda}$, which is one among the many curvature invariants in 4D Riemann spaces [121, 122]. In EF coordinates $K$ takes the form,
$K=\frac{1}{r^{4}M^{6}}\big{[}M^{\prime 2}(8r^{2}N^{2}+r^{4}N^{\prime 2})-2r^{2}MM^{\prime}N^{\prime}(4N+r^{2}N^{\prime\prime})+M^{2}(4(M^{2}-N)^{2}+4r^{2}N^{\prime 2}+r^{4}(N^{\prime\prime})^{2})\big{]}$ (72)
@@
The plots below show the Kretschmann scalar for some special cases of the horizon and asymptotic four potentials considered in figure 6 and 7 along with that for the RN case. These follow directly from the numerical code, and the same initial values are used.
(a) A plot of the Kretschmann Scalar $K=R^{\mu\nu\kappa\lambda}R_{\mu\nu\kappa\lambda}$ for the case of plot 6(d). Note that the curvature near $x=1$ is actually finite and converges to zero as the function $N$ does (not shown). We chose the vertical range to exhibit the interesting bump feature.
(b) A plot of the Kretschmann Scalar for the case of plot 7(c), showing a divergence as $x\rightarrow 1.90$.
@@
Our aim here is not a complete study of curvature invariants. However, we close with an interesting analytical result that can be used for to show interesting feature of the hypergeometric potentials versus others. Consider the Ricci scalar $R$ obtained by tracing the field equations (34). Due to the traceless nature of the field strength portion of the energy-momentum tensor, we find a general expression in terms of the potential only. For any choice of $V(X)$, where $X=B^{\mu}B_{\mu}$, we find a remarkably simple result,
@@
Figure 9: A plot of the (scaled) Ricci scalar for nonzero potential choices. For the Einstein-Maxwell case $R=0$. For the massive Proca potential and the quadratic potential, the Ricci scalar diverges for decreasing $X=B^{\alpha}B_{\alpha}/\Lambda^{2}$. For the hypergeometric potentials $R$ is bounded.
V.2 Asymptotic non-analytic solution
@@
First we linearize the equations (71) around asymptotic values (values at $u=0$).
@@
where the asymptotic values are indicated with a subscript 00, and $\epsilon$ is a small parameter to indicate a perurbative expansion. The argument of $V$ and $V^{\prime}$ is also expanded around an asymptotic value $-{\tilde{\beta}}=-f_{0}^{2}/N_{0}$, that minimizes the potential so that $V^{\prime}(-{\tilde{\beta}})=0$. Note that ${\tilde{\beta}}=\beta/\Lambda^{2}$, where $\beta$ was defined in Sec. IV.4. We also take the minimum value of the potential $\bar{V}$ to be zero. Upon inserting the expansion (74) in (71), and keeping first order in $\epsilon$ terms, we obtain the following linear equations for the perturbations:
@@
$u\,\delta N^{\prime}$ $=\delta N-2\sqrt{N_{0}}\delta M.$
@@
Solutions to these equations can be using standard methods by solving the coupled system (75). We find the following general solutions, neglecting extra constant terms,
@@
$\delta N$ =±$=\pm k\sqrt{{\tilde{\beta}}N_{0}}\left(c_{1}ue^{\sqrt{\rho}/u}+\frac{c_{2}}{2\sqrt{\rho}}ue^{-\sqrt{\rho}/u}\right)+\frac{2d_{2}}{2-k{\tilde{\beta}}}u,$
$\delta M$ =±$=\pm\frac{1}{2}k\sqrt{{\tilde{\beta}}\rho}\left(c_{1}ue^{\sqrt{\rho}/u}-\frac{c_{2}}{2\sqrt{\rho}}ue^{-\sqrt{\rho}/u}\right),$
@@
where $\rho=(k{\tilde{\beta}}-2)2{\tilde{\beta}}{\tilde{V}}^{\prime\prime}_{0}$ and $c_{1}$, $c_{2}$, and $c_{3}$ are constants. Note the interesting dependence on the minimum value of the vector field (scaled) ${\tilde{\beta}}$. We can have both oscillatory behavior or growing/decaying exponential behavior, similar to [98].
@@
We examine the full numerical solution “seeded” by these linearized approximations in order to determine how solutions that are not analytic at $u=0$ evolve with $u$. We confine ourselves to two examples, for brevity. One is the hypergeometric potential with $n=3$ and other is the quadratic potential. In figure 10, are two cases where ${\tilde{\beta}}=3$ and ${\tilde{\beta}}=1$, so that we capture damped behavior versus oscillatory. For the former case we eliminate the terms growing with small $u$, with the positive exponential argument. We use the same numerical scheme as described in section V.1.2, and input the initial values of $f$, $f^{\prime}$, $N$ and $M$ at small u values $u=1/20$.
@@
(a) The quadratic potential case with initial data from (75). This shows initial oscillatory behavior followed by a naked singularity.
@@
(b) Zoomed in portion showing the contrast of the numerical solutions with the solution from (75).
(c) The hypergeometric potential $M(3,2,X)$ case.
@@
(d) The hypergeometric potential $M(3,2,X)$ case again; zoomed-in portion of the left plot showing the contrast of the numerical solutions with the perturbation solution from (76).
@@
Figure 10: Numerical solutions using the non-analytic expansion (75) to generate initial values near $u=0$. The results (76) are labeled with subscript $na$. The solid curves are the numerical code, the dashed lines are the RN case, and the dotted lines show the approximate solutions (76). We use the values $N_{0}=1=M_{0}$, with ${\tilde{\beta}}=1$ (Quadratic) and ${\tilde{\beta}}=3$ (hypergeometric). We choose the constants to be “small” $c_{1}\sim 1/50$ for the quadratic case, and $c_{1}\sim 1/250$ for the Hypergeometric case, to make the perturbation solution approximately valid. The connection with conventional Schwarzschild coordinates is $g_{tt}=A=-N$
@@
The plots 10(a) and 10(c) show the overall behavior while 10(b) and 10(d) show behavior near $u=0$. Both cases show agreement between the numerical solution and the approximation (75) for small $u$. In the case of the quadratic potential, the numerical solution clearly shows the same, small $u$, oscillatory behavior. For the hypergeometric case, the approximate solution is not oscillatory because $\beta=3$ so $\rho>0$, but it tracks the numerical solution closely for small $u$. However, it appears both solutions have diverging $N$ as $u\rightarrow\infty$, indicating a naked singularity.
In this subsection, we employ numerical integration to study the trajectories of particles in free-fall in the spherically symmetric spacetime.

### v2

@@ 1 $a_{f1}$ $72e^{4}{\tilde{g}}a_{f1}^{2}/\beta_{3}^{2}$ $\beta_{3}/6e^{2}$
Now we consider the first hypergeometric potential $M(3,2,z)$. The behavior indicated by the sample in Fig. 6c suggests a regular solution for the functions $f$, $M$, and $N$ until one reaches $x=4.38$, where the numerical code fails (for the given initial conditions). We have analyzed this region (near $x=4.38$) with calculation of curvature scalars and we find disagreement (see the curvature section below). For example, the Ricci scalar we find must be finite at $x=4.38$, in disagreement with the numerical calculations. This behavior is consistent when one varies the initial values of $f^{\prime}$ and the coupling $\tilde{g}$}; different choices merely move the stiff point to larger or shorter $x$. Thus it remains an open question to decipher the behavior at these points, whether they are coordinate singularities or physical ones.
@@
For the case of $M(4,2,z)$, the results of the sample in Fig. 6d show an interesting pulse-like behavior similar to the quadratic potential case. Around $x=25$, the function $N$ hits zero again, indicated another possible horizon. Indeed, calculations of the curvature show that it is finite at $x=25$. Thus we can say that this case contains RN-like behavior with two horizons. The pulse-like behavior persists when one changes the horizon values of $f^{\prime}(1)$ and $\tilde{g}$}, as discussed below in Sec. V.1.3.
@@
Not shown is the case for $M(5,2,z)$, for which we find that the behavior is similar to the case $M(4,2,z)$. The calculation of the curvature also confirms that when $N$ approaches zero, the curvature is finite as indicated from curvature invariants. For both $n=4,5$ hypergeometric potentials it remains an open problem to map out the spacetime beyond this value of $x$ for which one gets horizon behavior.
@@ V.1.2 Asymptotic region
We then solve order by order in $u$. Note that this implies that the coefficients in the expansion must take values to cancel terms with inverse powers of $u$, starting with $1/u^{4}$. We include tables with the first three terms in the series for the different choices of potential.
@@ 1 $c_{f1}$ 0 $c_{N1}$ $c_{f1}$ 0 $2c_{f0}c_{f1}$
0 $c_{f0}$ $c_{f0}/\sqrt{3}$ $c_{f0}^{2}/3$ $c_{f0}$ $c_{f0}/\sqrt{2}$ $c_{f0}^{2}/2$
1 $c_{f1}$ 0 $\tfrac{2}{3}c_{f0}c_{f1}$ $c_{f1}$ 0 $c_{f0}c_{f1}$
2 $c_{f1}^{2}/c_{f0}$ $\frac{\sqrt{3}kc_{f1}^{2}}{2c_{f0}}$ 0 $c_{f1}^{2}/c_{f0}$ $kc_{f1}^{2}/(\sqrt{2}c_{f0})$ 0
@@ 1 $c_{f1}$ 0 $\tfrac{2}{3}c_{f0}c_{f1}$ $c_{f1}$ 0 $c_{f0}c_{f1}$
Tables 4 and 5 display some general features for how the functions behave in the asymptotic region. Firstly, the initial data needed to determine the behavior rely only on $f$ and $df/du$ at $u=0$. For the non-RN cases of the quadratic and hypergeometric potentials, the signs of $c_{f0}$ and $c_{f1}$ are critical to the behavior of the solutions for increasing $u$, and thus decreasing $x$. Regarding the function $N$, if $c_{f1}c_{f0}>0$ then $N$ will be increasing up to order $u^{3}$, and the order $u^{2}$ term vanishes. This behavior resembles that of a naked singularity. Indeed this is borne out by studying numerical solutions. On the other hand if $c_{f1}c_{f0}<0$, then $N$ will start off decreasing. For the latter condition, this is a good starting constraint on the initial values in order to try to find solutions that approach a horizon, like the conventional RN case.
The function $M$ has a different behavior. The $O(u)$ term vanishes for all potentials. The second and third order terms can be positive or negative depending on choices of $c_{f0}$ and $c_{f1}$. This was to be expected given the $M^{\prime}\sim M^{3}$ structure in equation (71).
@@
We can also confirm the Bekenstein result for the massive vector field, as the exact solution for this case emerges from the asymptotic expansion. Specifically, the solutions for the massive potential require that $c_{f0}=0=c_{f1}$, so this case reduces to the Schwarzschild solution with zero vector field.
@@
(a) The quadratic potential case with asymptotic values for the functions of order unity. The initial values are taken to be f(0)=1f(0)=-1, $f^{\prime}(0)=1/2$ and we set ${\tilde{\lambda}}=1/5$.
(b) A second case of the quadratic potential with initial values of f(0)=1f(0)=-1, $f^{\prime}(0)=-1/5$, and ${\tilde{\lambda}}=1$. This behavior is consistent with a naked singularity.
(c) The hypergeometric potential $M(3,2,z)$ case. The asymptotic values used are $f(0)=-\sqrt{3}$ and $f^{\prime}(0)=1/\sqrt{3}$ while ${\tilde{g}}=1$ and $k=1/2$.
(d) The hypergeometric potential $M(4,2,z)$ case. The initial values are $f(0)=-\sqrt{2}$ and $f^{\prime}(0)=1/\sqrt{2}$, while $k=2/3$ and ${\tilde{g}}=1$.
Figure 7: Numerical solutions using the asymptotic expansion (70) to generate initial values near $u=0$. The analytical results with the series expansion are labeled with subscript AA. The dashed lines are the RN case while the dotted lines show the analytical series from (70); the latter includes terms up to $u^{6}-u^{7}$. The connection with conventional Schwarzschild coordinates is $g_{tt}=A=-N$.
@@
For the quadratic potential, shown in Figs. 7a and 7b, two choices of initial values are made, one which yields decreasing behavior of $N$, the other increasing behavior. For the former, while $N$ appears headed to zero, $M$ increases without limit, in what appears to be a singular or stiff point in the numerical simulation. From curvature analysis, the point $u\approx 1.002$ appears to be a physical singularity of a different type from conventional ones (because $N\rightarrow 0$ and $M\rightarrow\infty$). We hesitate to call it a naked singularity because it may be partially hidden since $N\rightarrow 0$. For the latter case in Fig. 7b, qualitatively different boundary conditions yield a clear naked singularity.
@@
For the $M(3,2,z)$ case, we display a plot for which the asymptotic values will yield $dN/du<0$ initially. As shown in Fig. 7c, the solution indeed appears to have $N$ reach 0. However, at this point the numerical code breaks down. Studies of curvature scalars show a smoothly increasing Kretschmann curvature scalar, revealing that there appears to be a singularity at $u\approx 1.90$. The smoothness gives an indicator of the reliability of the numerical solver [120]. This is similar to the case encountered for the quadratic potential.
For the $M(4,2,z)$ case, we again display a plot for which the asymptotic values will yield $dN/du<0$ initially. As shown in Fig. 7d, the solution appears to have $N$ reach 0. At this point, the numerical code breaks down; the curvature scalars show a smoothly increasing Kretschmann scalar, again indicating a singularity at $u\approx 1.12$.
As in the horizon expansion case in Section V.1.1, it would be of interest to further study these cases over a range of values for the initial data. We leave this for future work.
@@ V.1.3 Curvature
Invariant quantities can show unequivocally if singularities exist in the solutions. We first examine the Kretschmann scalar $K=R_{\mu\nu\kappa\lambda}R^{\mu\nu\kappa\lambda}$, which is one among the many curvature invariants in four-dimensional Riemann spaces [121, 122]. In EF coordinates $K$ takes the form,
$K$ $=\frac{1}{r^{4}M^{6}}\big{[}M^{\prime 2}(8r^{2}N^{2}+r^{4}N^{\prime 2})-2r^{2}MM^{\prime}N^{\prime}(4N$ (72)
$\phantom{=}+r^{2}N^{\prime\prime})+M^{2}(4(M^{2}-N)^{2}+4r^{2}N^{\prime 2}+r^{4}(N^{\prime\prime})^{2})\big{]}$
@@
The plots below show the Kretschmann scalar for some special cases of the horizon and asymptotic expansions considered in Figs. 6 and 7, along with that for the RN case. These follow directly from the numerical code, and the same initial values are used. Note that for the Kretschmann scalar in Fig. 8a, the values of the potential strength $\tilde{g}$}, the gravity coupling strength $k$, and the horizon value of $f^{\prime}(1)$ are varied. This plot shows the persistence of a bump feature indicating the rapid change of the metric functions shown in 6d.
(a) A plot of the Kretschmann scalar $K=R^{\mu\nu\kappa\lambda}R_{\mu\nu\kappa\lambda}$ for the case of plot 6d, with small parameter variations. The curvature near $x=1$ is actually finite (not shown). We chose the vertical range to exhibit the interesting bump feature. The curves are distinguished by the values of $({\tilde{g}},k,f^{\prime}(1))$.
(b) A plot of the Kretschmann Scalar for the case of plot 7c, showing a divergence as $x\rightarrow 1.90$. This point is where the solution for NN tends to zero.
@@
Our aim here is not a complete study of curvature invariants. However, we close with an interesting analytical result that can be used to show an interesting feature of the hypergeometric potentials versus others. Consider the Ricci scalar $R$ obtained by tracing the field equations (34). Due to the traceless nature of the field strength portion of the energy-momentum tensor, we find a general expression in terms of the potential only. For any choice of $V(X)$, where $X=B^{\mu}B_{\mu}$, we find a remarkably simple result,
@@
Figure 9: Plot of the (scaled) Ricci scalar for nonzero potential choices. For the Einstein-Maxwell case $R=0$. For the massive Proca potential and the quadratic potential, the Ricci scalar diverges for decreasing $X/\Lambda^{2}$. For the hypergeometric potentials $R$ is bounded.
V.2 Asymptotic nonanalytic solution
@@
First we linearize the equations (71) around asymptotic values (values at $u=0$),
@@
where the asymptotic values are indicated with a subscript 0, and $\epsilon$ is a small parameter to indicate a perturbative expansion. The argument of $V$ and $V^{\prime}$ is also expanded around an asymptotic value ${\tilde{\beta}}=-f_{0}^{2}/N_{0}$, that minimizes the potential so that $V^{\prime}(-{\tilde{\beta}})=0$. Note that ${\tilde{\beta}}=\beta/\Lambda^{2}$, where $\beta$ was defined in Sec. IV.4. We also take the minimum value of the potential $\bar{V}$ to be zero. Upon inserting the expansions in equations (74) into (71), and keeping first order in $\epsilon$ terms, we obtain the following linear equations for the perturbations:
@@
$u\,\delta N^{\prime}$ $=\delta N-2\sqrt{N_{0}}\delta M$
@@
Solutions to these equations can be found using standard methods by solving the coupled system (75). We find the following general solutions, neglecting extra constant terms,
@@
$\delta N$ $=\pm k\sqrt{{\tilde{\beta}}N_{0}}\left(c_{1}ue^{\sqrt{\rho}/u}+\frac{c_{2}}{2\sqrt{\rho}}ue^{-\sqrt{\rho}/u}\right)$
$\phantom{=}+\frac{2d_{2}}{2-k{\tilde{\beta}}}u,$
$\delta M$ $=\pm\frac{1}{2}k\sqrt{{\tilde{\beta}}\rho}\left(c_{1}e^{\sqrt{\rho}/u}-\frac{c_{2}}{2\sqrt{\rho}}e^{-\sqrt{\rho}/u}\right),$
@@
where $\rho=(k{\tilde{\beta}}-2)2{\tilde{\beta}}{\tilde{V}}^{\prime\prime}_{0}$ and $c_{1}$, $c_{2}$, and $d_{2}$ are constants. Note the interesting dependence on the minimum value of the vector field (scaled) $\tilde{\beta}$}. We can have both oscillatory behavior or growing/decaying exponential behavior, similar to Ref. [99].
@@
We examine the full numerical solution “seeded” by these linearized approximations in order to determine how solutions that are not analytic at $u=0$ evolve with $u$. We confine ourselves to two examples, for brevity, and we set the constants $\tilde{\lambda}$} and $\tilde{g}$} to unity. One is the hypergeometric potential with $n=3$ and other is the quadratic potential. In Fig. 10, are two cases where ${\tilde{\beta}}=3$ and ${\tilde{\beta}}=1$, so that we capture damped behavior versus oscillatory. For the former case we eliminate the terms growing with small $u$, with the positive exponential argument. For the latter case, $c_{2}$ is related to $c_{1}$ such that the solutions are real-valued. We use the same numerical scheme as described in Section V.1.2, and input the initial values of $f$, $f^{\prime}$, $N$ and $M$ at small u values $u=1/20$.
@@
(a) The quadratic potential case with initial data from (75). This shows initial oscillatory behavior followed by a naked singularity
@@
(b) Close-up portion showing the contrast of the numerical solutions with the solution from (75).
(c) The hypergeometric potential $M(3,2,z)$ case.
@@
(d) The hypergeometric potential $M(3,2,z)$ case again; close-up portion of the left plot showing the contrast of the numerical solutions with the perturbation solution from (76).
@@
Figure 10: Numerical solutions using the nonanalytic expansion (75) to generate initial values near $u=0$. The results (76) are labeled with subscript nana. The solid curves are the numerical code, the dashed lines are the RN case, and the dotted lines show the approximate solutions (76). We use the values $N_{0}=1=M_{0}$, with ${\tilde{\beta}}=1$ (quadratic) and ${\tilde{\beta}}=3$ (hypergeometric). We choose the constants to be “small” $c_{1}\sim 1/50$ for the quadratic case, and $c_{2}\sim 1/250$ for the hypergeometric case, to make the perturbation solution approximately valid. The connection with conventional Schwarzschild coordinates is $g_{tt}=A=-N$.
@@
The plots 10a and 10c show the overall behavior while 10b and 10d show behavior near $u=0$. Both cases show agreement between the numerical solution and the approximation (75) for small $u$. In the case of the quadratic potential, the numerical solution clearly shows the same, small $u$, oscillatory behavior. For the hypergeometric case, the approximate solution is not oscillatory because $\beta=3$ so $\rho>0$, but it tracks the numerical solution closely for small $u$. However, it appears both solutions have diverging $N$ as $u\rightarrow\infty$, indicating a naked singularity.